# A Cartesian Cut-Cell Two-Fluid Method for Two-Phase Diffusion Problems

### v1

@@ 5.2 Diphasic validation
5.2.1 One-dimensional diphasic unsteady diffusion with homothetic jump
We consider a one-dimensional two-phase parabolic diffusion problem as a sanity check of the jump enforcement in a minimal setting. The domain is the interval $\Omega=[0,L]$ split by a fixed interface at $x=x_{\mathrm{int}}$ into
@@ 5.2.1 One-dimensional diphasic unsteady diffusion with homothetic jump
where $\lambda$ is prescribed.
Introduce the similarity variables
@@
and define the constant
$A=-\frac{\lambda}{1+\lambda},$ (66)
An exact solution satisfying (64)-(65) for all $t>0$ is given by
@@
The outer boundary conditions are then prescribed as time-dependent Dirichlet condition taken from (67)-(68):
@@
Table 6 reports $L^{2}$ errors at final time $t_{f}$ over regular, cut and all active cells. Pairwise convergence orders are also provided.
@@
In addition to the stiff case $\lambda=100$ reported below, we also consider a small sweep of homothetic jumps (e.g. $\lambda\in\{0.1,1,10,100\}$) in order to visualize the interface enforcement. As $\lambda$ varies, the analytical solution (67)-(68) transitions from a weak to a strong discontinuity in $\phi$ across $x_{\mathrm{int}}$, while the flux continuity constraint in (65) remains unchanged. This makes the test particularly sensitive to the discrete coupling: the method must simultaneously reproduce the prescribed jump amplitude and maintain consistent normal gradients on both sides of an interface that cuts a control volume.
Figure 15: One-dimensional diphasic unsteady diffusion with homothetic jump: numerical solution (markers) and analytical profiles (lines) at the final time $t_{f}$ for several jump ratios $\lambda$ (e.g. $\lambda=0.1,1,10,100$). The vertical dashed line indicates the interface position $x_{\mathrm{int}}$ (chosen inside a cut cell : $x_{int}=4.0+1e^{-2}$). The correct jump amplitude $\phi^{+}(x_{\mathrm{int}},t_{f})=\lambda\,\phi^{-}(x_{\mathrm{int}},t_{f})$ is recovered while preserving flux continuity across the interface.
@@ 5.2.2 Two-phase unsteady diffusion with a circular interface
The computational domain is the square cavity $\Omega=[0,L]^{2},\qquad L=8,$ containing a circular interface $\Gamma$ of radius $R_{0}=2$ centered at $\mathbf{x}_{c}=(4,4)$. The phases are defined by $\Omega^{+}=\{\mathbf{x}\in\Omega:\ r<R_{0}\},\Omega^{-}=\{\mathbf{x}\in\Omega:\ r>R_{0}\},r=\|\mathbf{x}-\mathbf{x}_{c}\|.$ In each phase we solve the unsteady diffusion equation
$c_{p}^{\pm}\,\frac{\partial\phi^{\pm}}{\partial t}=\nabla\cdot\bigl(K^{\pm}\,\nabla\phi^{\pm}\bigr)\qquad\text{in }\Omega^{\pm},$ (69)
with homogeneous Neumann conditions on the outer box boundary, $\partial_{n}\phi^{\pm}=0\qquad\text{on }\partial\Omega,$ and the interfacial matching conditions on $\Gamma$,
$\llbracket\phi\rrbracket=0,\qquad\llbracket K\,\nabla\phi\cdot\mathbf{n}\rrbracket=0,\qquad\text{on }\Gamma.$ (70)
The initial data are chosen as
$\phi^{+}(\mathbf{x},0)=\phi_{0},\qquad\phi^{-}(\mathbf{x},0)=0.$ (71)
@@
Table 7 reports the $L^{2}$ errors at the final time $t_{f}$ over regular cells, cut cells and all active cells in both phases. Several features are worth noting.
@@
The temporal evolution of the interfacial diffusive flux provides a stringent check of the two-phase coupling. Table 8 shows that the final flux value $Q^{+}(t_{f})$ converges monotonically toward its reference value as the mesh is refined, with the relative error decreasing from $3.56\times 10^{-2}$ at $h=0.5$ to $1.42\times 10^{-3}$ at $h=0.0625$. The corresponding rates are close to first-to-second order over this range, indicating that both the enforcement of the interfacial condition (70) and the reconstruction of normal gradients in cut cells remain accurate. Figure 16 further illustrates that the normalized interfacial flux history collapses under refinement: transient discrepancies on coarse grids diminish confirming that the method captures the time-dependent exchange across the embedded interface.
@@
5.2.3 Diphasic unsteady diffusion across a sphere
We conclude the validation with a fully three-dimensional diphasic test in a spherical geometry, for which a closed analytical reference is available in integral form (brown_diffusion_1965). This benchmark simultaneously checks the two-fluid coupling across a curved interface through scalar continuity and flux continuity and the robustness of the 3D moment integration required by the cut-cell discretization.
Let $\mathbf{x}_{c}=(2,2,2)$ and $R_{0}=1$. The interface is the sphere $\Gamma=\{\mathbf{x}:\ r=R_{0}\},r=\|\mathbf{x}-\mathbf{x}_{c}\|,$ splitting $\Omega_{\rm box}=[0,4]^{3}$ into the inner phase $\Omega^{-}=\{r<R_{0}\}$ and the outer phase $\Omega^{+}=\{r>R_{0}\}$. We solve
$\partial_{t}\phi^{\pm}-D^{\pm}\Delta\phi^{\pm}=0\qquad\text{in }\Omega^{\pm},$ (74)
with diffusivities $D^{-}=D_{g}$ and $D^{+}=D_{l}$. Across $\Gamma$, we impose scalar continuity and normal-flux continuity:
$\phi^{+}(R_{0},t)=\phi^{-}(R_{0},t),\qquad D_{l}\,\partial_{r}\phi^{+}(R_{0},t)=D_{g}\,\partial_{r}\phi^{-}(R_{0},t),$ (75)
where $\partial_{r}$ denotes the radial derivative. The analytical solution corresponds to the classical configuration of an initially uniform inner concentration and vanishing outer concentration,
$\phi^{-}(r,0)=c_{0},\quad 0\leq r<R_{0},\qquad\phi^{+}(r,0)=0,\quad r>R_{0},$ (76)
@@ 5.2.3 Diphasic unsteady diffusion across a sphere
The solution is extracted from brown_diffusion_1965. The solution is radially symmetric, $\phi^{\pm}(\mathbf{x},t)=\Phi^{\pm}(r,t)$. Following brown_diffusion_1965, define the auxiliary parameters
$\sigma=\sqrt{\frac{D_{l}}{D_{g}}},\qquad Q=\frac{D_{l}}{D_{g}}\,\sigma,\qquad L=\frac{D_{l}-D_{g}}{D_{g}},$ (77)
@@
$\Phi^{-}(r,t)=\frac{2Q\alpha c_{0}}{\pi r}\int_{0}^{\infty}\frac{\bigl(\sin u-u\cos u\bigr)\,\sin\!\left(u\,\frac{r}{R_{0}}\right)}{\mathcal{D}(u)}\;\exp\!\left(-D_{g}\,\frac{u^{2}}{R_{0}^{2}}\,t\right)\,{\rm d}u.$ (79)
@@
$\mathcal{F}(u,r)=\bigl(u\cos u+L\sin u\bigr)\,\sin\!\left(u\,\frac{r-R_{0}}{\sigma R_{0}}\right)+Q\,u\sin u\,\cos\!\left(u\,\frac{r-R_{0}}{\sigma R_{0}}\right),$ (80)
@@
$\Phi^{+}(r,t)=\frac{2\alpha c_{0}}{\pi r}\int_{0}^{\infty}\frac{\bigl(\sin u-u\cos u\bigr)\,\mathcal{F}(u,r)}{u\,\mathcal{D}(u)}\;\exp\!\left(-D_{g}\,\frac{u^{2}}{R_{0}^{2}}\,t\right)\,{\rm d}u.$ (81)
In practice, the integrals (79)-(81) are evaluated by adaptive quadrature and truncated at a large cutoff. Time integration uses a midpoint time scheme, with a time step $\Delta t=0.25\min(\Delta x,\Delta y,\Delta z)$.
Table 9 reports the $L^{2}$ errors at the final time $t_{f}$ over all active cells, regular cells and cut cells. The systematic decay of these three measures under mesh refinement confirms that the two-fluid diphasic coupling remains consistent in three dimensions. This benchmark is also a stringent stress test for the geometric layer: a spherical interface generates a broad spectrum of cut-cell patterns, including very small trimmed control volumes, strongly skewed face apertures and highly curved interface patches. The convergence observed in Table 9 therefore indicates that the three-dimensional moment integration (phase volumes, face apertures and interface measures/centroids) is sufficiently accurate and robust to sustain stable discrete flux balances and sharp enforcement of the interface conditions (75). The pairwise orders are not uniform across all refinement steps. For coarse meshes, the sphere is under-resolved and changes in cut-cell topology can temporarily shift the relative weight of bulk and interface contributions, leading to pre-asymptotic behavior. As the resolution increases, the regular- cell and global errors display a clear rapid decay, whereas the cut-cell error may exhibit apparent superconvergence. Overall, the finest refinements exhibit superlinear convergence in the global measures.
@@
Table 9: Diphasic unsteady diffusion in a sphere (Brown solution): $L^{2}$ errors at final time $t_{f}$ over all, regular and cut cells, with pairwise convergence orders. The row “fit” reports a least-squares slope of $\log(\|e\|)$ versus $\log(h)$.
@@ 6 Conclusions and outlook
This work introduced a conservative Cartesian cut-cell finite-volume method for sharp-interface diffusion problems in static embedded geometries, in both monophasic and diphasic configurations. The formulation relies on a reduced geometric description: phase-restricted volumes, face apertures and a small set of low-order geometric moments that are sufficient to assemble compact stencils for discrete divergence and gradient operators. Boundary conditions (Dirichlet, Neumann and Robin) and interface jumps (normal diffusive flux and value) are enforced strongly in the algebraic system. A distinctive feature of the two-fluid approach is that interfacial unknowns are explicit degrees of freedom, providing direct access to interface values without post-processing or interpolation.
The numerical validation in static domains demonstrates accuracy, conservation, and robustness across a range of embedded geometries. Monophasic elliptic tests on non-convex domains recover near second-order convergence in $L^{2}$ norm, with cut-cell errors converging at comparable rates. The discrete solution remains within the expected physical range, with no overshoot or undershoot observed as the grid is refined. For Robin boundary condition, the method maintains stable convergence in both $L^{2}$ and $H^{1}$ norms and preserves a faithful reconstruction of boundary unknowns. In three dimensions, the method extends without modification and remains effective even at very low resolutions, when the embedded body is represented by only a few cells. In the Neumann limit, the discrete mass is preserved to machine precision, highlighting exact flux conservation at the fully discrete level. Diphasic validations confirm that sharp coupling across interfaces does not compromise robustness or accuracy. The method exhibits near second-order convergence in global norms and appropriate behavior in cut cells, while the interfacial flux converges rapidly under refinement. Even with a strong homothetic jump, the coupling remains robust. Both the jump enforcement and the flux continuity remain well behaved as the interface cuts control volumes at arbitrary locations. Overall, the results show that the proposed method provides a reliable and geometrically flexible building block for interface-resolved diffusion with compact conservative operators constructed from a reduced set of moments, direct discrete enforcement of boundary and interface conditions, robust behavior in severely under-resolved regimes and consistent accuracy in both mono- and two-phase configurations.

### v2

@@ 5.2 Two-phase validation
5.2.1 One-dimensional two-phase unsteady diffusion with homothetic jump
We consider a one-dimensional two-phase parabolic diffusion problem as a check of the jump enforcement in a minimal setting. The domain is the interval $\Omega=[0,L]$ split by a fixed interface at $x=x_{\mathrm{int}}$ into
@@ 5.2.1 One-dimensional two-phase unsteady diffusion with homothetic jump
where $\lambda$ is prescribed. A self-similar solution is obtained by introducing the similarity variables,
@@
and defining the constant
$A=-\frac{\lambda}{1+\lambda}.$ (66)
The exact solution to the problem Eqs. (64)-(65) for all $t>0$ is then given by
@@
The outer boundary conditions are then prescribed as time-dependent Dirichlet condition taken from Eqs. (67)-(68):
@@
Tab. 6 reports $L^{2}$ errors at final time $t_{f}$ over regular, cut and all active cells. Pairwise convergence orders are also provided.
@@
In addition to the stiff case $\lambda=100$ reported below, we also consider a small sweep of homothetic jumps (e.g. $\lambda\in\{0.1,1,10,100\}$) in order to visualize the interface enforcement. As $\lambda$ varies, the analytical solution Eqs. (67)-(68) transitions from a weak to a strong discontinuity in $\phi$ across $x_{\mathrm{int}}$, while the flux continuity constraint in Eq. (65) remains unchanged. This makes the test particularly sensitive to the discrete coupling: the method must simultaneously reproduce the prescribed jump amplitude and maintain consistent normal gradients on both sides of an interface that cuts a control volume.
Figure 14: One-dimensional two-phase unsteady diffusion with homothetic jump: numerical solution (markers) and analytical profiles (lines) at the final time $t_{f}$ for several jump ratios $\lambda$ (e.g. $\lambda=0.1,1,10,100$). The vertical dashed line indicates the interface position $x_{\mathrm{int}}$ (chosen inside a cut cell : $x_{int}=4.0+1e^{-2}$). The correct jump amplitude $\phi^{+}(x_{\mathrm{int}},t_{f})=\lambda\,\phi^{-}(x_{\mathrm{int}},t_{f})$ is recovered while preserving flux continuity across the interface.
@@ 5.2.2 Two-phase unsteady diffusion with a circular interface
The computational domain is the square cavity $\Omega=[0,L]^{2},L=8,$ containing a circular interface $\Gamma$ of radius $R_{0}=2$ centered at $\mathbf{x}_{c}=(4,4)$. The phases are defined by
$\Omega^{+}=\{\mathbf{x}\in\Omega:\ r<R_{0}\},\quad\Omega^{-}=\{\mathbf{x}\in\Omega:\ r>R_{0}\},\quad r=\|\mathbf{x}-\mathbf{x}_{c}\|.$
In each phase, we solve the unsteady diffusion equation with homogeneous Neumann conditions on the outer boundary $\partial\Omega$. At the interface, we consider the continuity case $\lambda=1.0$, i.e. both the field and the normal diffusive flux are continuous across $\Gamma$. The initial data are chosen as $\phi^{+}(\mathbf{x},0)=\phi_{0}$ and $\phi^{-}(\mathbf{x},0)=0$.
@@
Tab. 7 reports the $L^{2}$ errors at the final time $t_{f}$ over regular cells, cut cells and all active cells in both phases. Several features are worth noting.
@@
The temporal evolution of the interfacial diffusive flux provides a stringent check of the two-phase coupling. Tab. 8 shows that the final flux value $Q^{+}(t_{f})$ converges monotonically toward its reference value as the mesh is refined, with the relative error decreasing from $3.56\times 10^{-2}$ at $h=0.5$ to $1.42\times 10^{-3}$ at $h=0.0625$. The corresponding rates are close second order over this range, indicating that both the enforcement of the interfacial condition and the reconstruction of normal gradients in cut cells remain accurate. Fig. 15 further illustrates that the normalized interfacial flux history collapses under refinement: transient discrepancies on coarse grids diminish confirming that the method captures the time-dependent exchange across the embedded interface.
@@
5.2.3 Two-phase unsteady diffusion across a sphere
We conclude the validation with a three-dimensional two-phase test in a spherical geometry, for which a closed analytical reference is available in integral form (brown_diffusion_1965). This benchmark simultaneously checks the two-fluid coupling across a curved interface through scalar continuity and flux continuity and the robustness of the 3D moment integration required by the cut-cell discretization.
Let $\mathbf{x}_{c}=(2,2,2)$ and $R_{0}=1$. The interface is the sphere $\Gamma=\{\mathbf{x}:\ r=R_{0}\},r=\|\mathbf{x}-\mathbf{x}_{c}\|,$ splitting $\Omega_{\rm box}=[0,4]^{3}$ into the inner phase $\Omega^{-}=\{r<R_{0}\}$ and the outer phase $\Omega^{+}=\{r>R_{0}\}$. We consider unsteady diffusion in the two phases $\Omega^{\pm}$, with constant diffusivities $D^{-}$ and $D^{+}$. Across the interface $\Gamma$ at $r=R_{0}$, we enforce continuity of the scalar field together with continuity of the normal diffusive flux.
The analytical solution corresponds to the classical configuration of an initially uniform inner concentration and vanishing outer concentration,
$\phi^{-}(r,0)=\phi_{0},\quad 0\leq r<R_{0},\qquad\phi^{+}(r,0)=0,\quad r>R_{0},$ (71)
@@ 5.2.3 Two-phase unsteady diffusion across a sphere
The solution is radially symmetric, $\phi^{\pm}(\mathbf{x},t)=\Phi^{\pm}(r,t)$. Following brown_diffusion_1965, we introduce the auxiliary parameters
$\sigma=\sqrt{D^{+}/D^{-}},\qquad Q=\frac{D^{+}}{D^{-}}\,\sigma,\qquad L=\frac{D^{+}-D^{-}}{D^{-}},$ (72)
@@
$\Phi^{-}(r,t)=\frac{2Q\alpha\phi_{0}}{\pi r}\int_{0}^{\infty}\frac{\bigl(\sin u-u\cos u\bigr)\,\sin\!\left(ur/R_{0}\right)}{\mathcal{D}(u)}\;\exp\!\left(-D^{-}u^{2}t/R_{0}^{2}\right)\,{\rm d}u.$ (74)
@@
$\mathcal{F}(u,r)=\bigl(u\cos u+L\sin u\bigr)\,\sin\!\left(u(r-R_{0})/\sigma R_{0}\right)+Q\,u\sin u\,\cos\!\left(u(r-R_{0})/\sigma R_{0}\right),$ (75)
@@
$\Phi^{+}(r,t)=\frac{2\alpha\phi_{0}}{\pi r}\int_{0}^{\infty}\frac{\bigl(\sin u-u\cos u\bigr)\,\mathcal{F}(u,r)}{u\,\mathcal{D}(u)}\;\exp\!\left(-D^{-}u^{2}t/R_{0}^{2}\right)\,{\rm d}u.$ (76)
In practice, the integrals Eqs. (74)-(76) are evaluated by adaptive quadrature and truncated at a large cutoff. Time integration uses a midpoint time scheme, with a time step $\Delta t=0.25\min(\Delta x,\Delta y,\Delta z)$.
Tab. 9 reports the $L^{2}$ errors at the final time $t_{f}$ over all active cells, regular cells and cut cells. The systematic decay of these three measures under mesh refinement confirms that the two-fluid two-phase coupling remains consistent in three dimensions. This benchmark is also a stringent stress test for the geometric layer: a spherical interface generates a broad spectrum of cut-cell patterns, including very small trimmed control volumes, strongly skewed face apertures and highly curved interface patches. The convergence observed in Tab. 9 therefore indicates that the three-dimensional moment integration (phase volumes, face apertures and interface measures/centroids) is sufficiently accurate and robust to sustain stable discrete flux balances and sharp enforcement of the interface conditions. The pairwise orders are not uniform across all refinement steps. For coarse meshes, the sphere is under-resolved and changes in cut-cell topology can temporarily shift the relative weight of bulk and interface contributions, leading to pre-asymptotic behavior. As the resolution increases, the regular- cell and global errors display a clear rapid decay, whereas the cut-cell error may exhibit apparent superconvergence.
@@
Table 9: Two-phase unsteady diffusion in a sphere : $L^{2}$ errors at final time $t_{f}$ over all, regular and cut cells, with pairwise convergence orders.
@@ 6 Conclusions and outlook
This work introduced a conservative Cartesian cut-cell finite-volume method for sharp-interface diffusion problems in static embedded geometries, in both single-phase and two-phase configurations. The formulation relies on a reduced geometric description: phase-restricted volumes, face apertures and a small set of low-order geometric moments that are sufficient to assemble compact stencils for discrete divergence and gradient operators. Boundary conditions (Dirichlet, Neumann and Robin) and interface jumps (normal diffusive flux and value) are enforced strongly in the algebraic system. A distinctive feature of the two-fluid approach is that interfacial unknowns are explicit degrees of freedom, providing direct access to interface values without post-processing or interpolation.
The numerical validation in static domains demonstrates accuracy, conservation, and robustness across a range of embedded geometries. Single-phase elliptic tests on non-convex domains recover near second-order convergence in $L^{2}$ norm, with cut-cell errors converging at comparable rates. The discrete solution remains within the expected physical range, with no overshoot or undershoot observed as the grid is refined. For Robin boundary condition, the method maintains robust convergence in both $L^{2}$ and $H^{1}$ norms and preserves a faithful reconstruction of boundary unknowns. In three dimensions, the method extends without modification and remains effective even at very low resolutions, when the embedded body is represented by only a few cells. In the Neumann limit, the discrete mass is preserved to machine precision, highlighting exact flux conservation at the discrete level. Two-phase validations confirm that sharp coupling across interfaces does not compromise robustness or accuracy. The method exhibits near second-order convergence in global norms and appropriate behavior in cut cells, while the interfacial flux converges rapidly under refinement. Even with a strong homothetic jump, the coupling remains robust. Both the jump enforcement and the flux continuity remain well behaved as the interface cuts control volumes at arbitrary locations. Overall, the results show that the proposed method provides a reliable and geometrically flexible building block for interface-resolved diffusion with compact conservative operators constructed from a reduced set of moments, direct discrete enforcement of boundary and interface conditions, robust behavior in severely under-resolved regimes and consistent accuracy in both mono- and two-phase configurations.